OpenStudy (anonymous):

calc help, integral [e^4 , e] 1/xlnx^(3)

OpenStudy (klimenkov):

\(\int\limits_{e}^{e^4}\frac{dx}{x \ln x^3}\) or \(\int\limits_{e}^{e^4}\frac{dx}{x \ln^3 x}\)

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

id use u substitution for that right?

OpenStudy (klimenkov):

Yes. Can you guess what substitution you need?

OpenStudy (anonymous):

lnx

OpenStudy (anonymous):

dx=dux

OpenStudy (klimenkov):

You are right. Don't forget to change limits while making substitution.

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (anonymous):

If I have allowance to ask something, I'd try to use \[u= ln(x^3)\] as a substitution, is that what you meant kingpin?

OpenStudy (anonymous):

I had left the ^3 in the equation. is that also going to be substituted?

OpenStudy (anonymous):

Because then you would obtain: \[ \frac{du}{dx}=\frac{1}{x^3} \cdot 3x^2 \\ \text{or} \\dx=\frac{x}{3}du \]

OpenStudy (anonymous):

then you have \[ \frac{1}{3} \int_a^b \frac{1}{x \cdot u } \cdot \frac{x}{3} du \]

OpenStudy (klimenkov):

\[u=\ln x^3, \ du=\frac{1}{x^3}\cdot 3x^2dx=\frac{3}xdx\]

OpenStudy (anonymous):

exactly @klimenkov

OpenStudy (klimenkov):

But there were limits for one function, and now you have another, so you will have new limits.

OpenStudy (anonymous):

one sec just working it out so i can see it myself

OpenStudy (anonymous):

the derivative of ln(x)^3 = 1/x^3 correct

OpenStudy (anonymous):

I dislike being a smarts but in this case it's needed. ln(x)^3 is not the same as ln(x^3)

OpenStudy (anonymous):

ohhhhhhhh, it would be 3(1/x)^2

OpenStudy (anonymous):

well yes, but in the equation it's \[ \ln(x^3) \] so the derivative is \[ \frac{1}{x^3} \cdot 3x^2 \]

OpenStudy (anonymous):

sorry if im being difficult i just learned this today

OpenStudy (anonymous):

you are doing fine, just showing empathy for the accurate way of typing it out, because the brackets can make a huge difference (-:

OpenStudy (anonymous):

basically you are using the chain rule then

OpenStudy (anonymous):

exactly!

OpenStudy (klimenkov):

Example: \[\int\limits_{1}^{e}\frac{\ln x}{x}\ dx\]Substitution: \(u=\ln x, \ du=\frac{dx}{x}\) New limits: \(a_1=u(a)=\ln 1=0, \ b_1=u(b)=\ln e=1\) Result: \[\int\limits_{1}^{e}\frac{\ln x}{x}\ dx=\int\limits_{0}^{1}u\ du\]

OpenStudy (anonymous):

alright so after finding the dx the new equation would be \[3 \int\limits_{e}^{e^4} du/xu\]

OpenStudy (anonymous):

which is the same as \[3\int\limits_{e}^{e^4} \left| u \right| /x\]?

OpenStudy (klimenkov):

Thats not right. You have to get rid of x.

OpenStudy (anonymous):

the x was canceled from before?

OpenStudy (klimenkov):

See my example and yours. I have no x and you have.

OpenStudy (klimenkov):

You can make the integral like in my example from yours problem by making something with logarithm. What do you have to do?

OpenStudy (anonymous):

i just have to evaluate the equation

OpenStudy (klimenkov):

Try to transform \(\ln x^3\).

OpenStudy (anonymous):

ok i just redid the whole problem what i cam up with is \[1/3\int\limits_{e}^{e^4} du \div u\]

OpenStudy (anonymous):

i know that the integral of du/u is \[\left| u \right|\]

OpenStudy (anonymous):

no, the integral of du/u is ln|u|

OpenStudy (anonymous):

yeah i just forgot to put the ln

OpenStudy (anonymous):

hehe ok, good then.

OpenStudy (anonymous):

if im using the e^4 and e it would look like (1/3) ln (ln(x^3)) - (1/3)ln(lnx^3))?

OpenStudy (anonymous):

well done, yes.

OpenStudy (anonymous):

and i can straight plug n chug in the calculator right?

OpenStudy (anonymous):

yes you can.